# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v1

@@ Appendix A Proofs
The inequality follows by Theorem 3.1 in [31]. $\Box$
The following lemma completes the proof of Theorem 15.
@@
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3. Then, for $x\in\mathcal{A}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
@@
$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{\|}_{\operatorname{op}}$
$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x)\|^{2q}_{\operatorname{op}}.$
@@
for any $z\in\mathbb{T}$. Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$. We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ [32, Chapter IX, Section 2].
In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$. Thus, we have
@@ Appendix B Details of Remark 10
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 9 from the perspective of $q$. The following lemma is by Brandlini and Travaglini [25].
@@
$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j,$
where $\hat{\cdot}$ is the Fourier transform, and the third equality is by the Poisson summation formula. $\Box$
To understand the convrgence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 9, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 [25, Theorem 1]. Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@ Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. [33, Proposition 4.5] showed an approximate representer theorem for RKHMs over von Neumann-algebras. Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
@@
Proof  For $f_{1},f_{2}\in\mathcal{M}_{k}$, let $\theta_{f_{1},f_{2}}:\mathcal{M}_{k}\to\mathcal{M}_{k}$ defined as $\theta_{f_{1},f_{2}}(v)=f_{1}\left\langle f_{2},v\right\rangle_{k}$ for $v\in\mathcal{M}_{k}$. Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, and let $\mathcal{B}=\overline{\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$. In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see [12, p8]). Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), it has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ [23, Theorem I.4.8]. For each $v\in\mathcal{N}_{k}$, we have
@@
Appendix D Experimental Details and Additional Results
@@
Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order. (Average value of results of five different runs. The error bar represents the standard deviation.)
D.1 Experiment with synthetic data
@@
D.2 Experiment with MNIST
We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.

### v3

@@ Appendix A Proofs
The inequality follows by Theorem 3.1 in Mohri et al. (2012). $\Box$
According to Hashimoto et al. (2023a), the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
@@
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Then, we have
Combining these results with the following lemma completes the proof of Theorem 16.
Lemma 24
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3. Then, for $x\in\mathcal{A}_{0}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
@@
$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(x_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(x_{i})^{q}\bigg{\|}_{\operatorname{op}}$
$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x_{i})\|^{2q}_{\operatorname{op}}.$
@@
for any $z\in\mathbb{T}$. Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}^{n\times n}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$. We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ (Conway, 2007, Chapter IX, Section 2).
In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$ and $\|v\|=1$. Thus, we have
@@ Appendix B Details of Remark 11
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 10 from the perspective of $q$. The following lemma is by Brandlini and Travaglini Brandolini & Travaglini (1997).
@@
$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}u_{j,t}(r)\mathrm{d}j$
$=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
$=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}nr\cdot m}\mathrm{d}r\mathrm{d}j,$
where $\hat{\cdot}$ is the Fourier transform, and the forth equality is by the Poisson summation formula. $\Box$
To understand the convergence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 10, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 (Brandolini & Travaglini, 1997, Theorem 1). Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
@@ Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. (Hashimoto et al., 2023b, Proposition 4.5) showed an approximate representer theorem for RKHMs over von Neumann-algebras. Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
@@
Proof  For $f_{1},f_{2}\in\mathcal{M}_{k}$, let $\theta_{f_{1},f_{2}}:\mathcal{M}_{k}\to\mathcal{M}_{k}$ defined as $\theta_{f_{1},f_{2}}(v)=f_{1}\left\langle f_{2},v\right\rangle_{k}$ for $v\in\mathcal{M}_{k}$. Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, let $\mathcal{B}_{0}={\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$, and let $\mathcal{B}=\overline{\mathcal{B}_{0}}$. In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see (Lance, 1995, p8)). Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), $\mathcal{B}_{0}$ has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ (Davidson, 1996, Theorem I.4.8). For each $v\in\mathcal{N}_{k}$, we have
@@
Appendix D Generalization to Other $C^{*}$-algebras
Setting $R_{n}$ and $S_{n}$ for more general $C^{*}$-algebras has been investigated. Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
Continuous functions on high-dimensional torus (Leimbach & van Suijlekom, 2024)
[…]
Continuous functions on the sphere (Rieffel, 2004)
Appendix E Experimental Details and Additional Results
@@
Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the regression task. (Average value of results of five different runs. The error bar represents the standard deviation.)
E.1 Experiment with synthetic data
@@
We also show additional results regarding the deep approach in Figure 4. The result is the average $\pm$ the standard deviation of three independent runs. We set different values of $n$ with the same deep setting considered in Subsection 7.1.
Figure 4: Test error of the regression task with deep approach with different $n$ and $L$.
E.2 Experiment with MNIST
We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. Figure 5 (a) shows the test error. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 5 (b) shows the result. The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
Figure 5: (a) Test error of the image recovering task with different values of $n$ and the kernel $k=\hat{k}_{n}^{\operatorname{prod},q}$. (Box plot of results of five independent runs with different training and test data.) (b) Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order for the image recovering task. (Average value of results of five different runs. The error bar represents the standard deviation.)